OpenStudy (anonymous):

400 millimeters x____=_____cm

OpenStudy (anonymous):

Is it your question : 400 ml . x = 1 cm Is it your question ? It is not complate !

OpenStudy (anonymous):

thats how the question looks on my paper. I'm so confused

OpenStudy (anonymous):

It should be have more number ( One number )

OpenStudy (anonymous):

Ok ! Look ! I say you thinks you can answer to your self . So just please answer !OK?!

OpenStudy (anonymous):

okkkkk

OpenStudy (anonymous):

Which one is the bigger ? 1 mil or 1 cm ?

OpenStudy (anonymous):

OK !!! No problem you should know this : 1 cm = 1000 ml

OpenStudy (unklerhaukus):

\[1{[\text{cm}] }={?[\text{mm}]}\\ \frac{[\text{cm}] }{?[\text{mm}]}=1\] \[400 [\text{mm}] =400 [\text{mm}] \times\frac{[\text{cm}] }{?[\text{mm}] } = 400 \color{red}{\cancel{\color{black}{[\text{mm}]}}} \times\frac1{?}\frac{[\text{cm}] }{\color{red}{{\cancel{\color{black}{[\text{mm}]}}}} } =\frac{400}?[\text{cm}]\]

OpenStudy (anonymous):

1 cm = 10mm impllies 400 mm =40(10mm) therefore 40 cm

OpenStudy (mathstudent55):

\(400 ~mm \times \dfrac{1~cm}{10~mm} = ~~~?~~~cm\)